OpenStudy (anonymous):

Your piggy bank has a total of 40 coins in it; some are dimes and some are nickels. If you have a total of $3.50, how many dimes do you have? Answer 30 10 40 20 Your piggy bank has a total of 40 coins in it; some are dimes and some are nickels. If you have a total of $3.50, how many dimes do you have? Answer 30 10 40 20 @Mathematics

hero (hero):

x + y = 40 .05x + .10y = 3.50 y = 40 - x y = (3.50 - .05x)/.10 y = y .10(40 - x) = 3.50 - .05x 4 - .1x = 3.50 - .05x 4 - 3.50 = .05x .50 = .05x x = .50/.05 = 10 y = 30 so she has 30 dimes